Use a numerical example to explain the difference between the rate of change and the level of change between two values.


1
Expert's answer
2021-06-10T10:58:54-0400

A level change is the difference in the level amount from one period to the next.

Rate of change is a percentage change that tells how fast an economic variable is rising or falling over a certain period of time.

The table below shows the numeric example.



From the above table;

Level of change is calculated by subtracting the previous year employment from the current year employment. For example, the level of change for the year 2009 is calculated as : 10,298.2-10,667.0 = -368.8

Rate of change is calculated using the formula

"[(\frac{t}{t-1})-1].100"

where, where t is equal to current year employment and t-1 is equal to the previous year.

The rate of change for 2008 is calculated as "[(\frac{10,298.2}{10,667.0})-1].100=-3.46".